Find the equation of the following line, write the equation in the slope intercept form. Identify the slope and y-intercept.
Through (1,4) and parallel to x – 2y = 6

Answer by tutmath(14)

Parallel lines will have the same slope.
Slope intercept form is
y = mx + c
where m is the slope and c is the y-intercept of the line
The line that satisfies the given equation:
x – 2y = 6
can be written as
y = x/2 - 3
Slope = 1/2
A line parallel to this line will also have slope = 1/2
Let us assume the equation of the line is
y = mx + c
Now we know m = 1/2 from above, also given that (1,4) is a point on the line so it satisfies the equation of the line.
Substituting,
4 = 1/2(1) + c
therefore, c = 7/2
So the equation of the line passing through (1,4) with slope 1/2 will be:
y = x/2 + 7/2
(tutmath@gmail.com)